OpenStudy (anonymous):

Please help me...Evaluate (-4)^-2

OpenStudy (anonymous):

= 1/(-4)^2 =1/(-4*-4) = 1/16

OpenStudy (jdoe0001):

keep in mind that \(\bf \Large a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1/(-4)^-2?

OpenStudy (isaiah.feynman):